OpenStudy (anonymous):

Approximate the logarithm using the properties of logarithms, given the values logb 2 ≈ 0.3562, logb 3 ≈ 0.5646, and logb 5 ≈ 0.8271. Round your result to four decimal places. logb 16/25

OpenStudy (anonymous):

log(16/25) = log(16) - log(25) = log(2^4) - log(5^2) = 4 log(2) - 2 log(5) you're the value of log(2) and log(5). Plug and chug

OpenStudy (jdoe0001):

\(\large log_b\left(\frac{16}{25}\right)?\quad or\quad log_{16}(25)?\)

OpenStudy (anonymous):

the first one

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

well... do we know what "b" is... is we're meant to get "b" as well as its logarithm?

OpenStudy (anonymous):

not sure, all i know is that the answer is a decimal

OpenStudy (jdoe0001):

did you try entering that in your calculator? does that show the same as the answer ?

OpenStudy (anonymous):

calculator is broken...

OpenStudy (jdoe0001):

https://www.desmos.com/calculator <---